OpenStudy (anonymous):

Calculus - I have a few of these problems answered, just checking my answers. 1. if f(x)+g(x)dx=8 [1,7] and f(x)dx=13 what does g(x) =?

OpenStudy (anonymous):

I got -5 as my answer because of the property of integrals

OpenStudy (anonymous):

You're correct!

OpenStudy (anonymous):

okay next one. 2. if F(x)=f(t)dt where f(t) = \[\int\limits_{1}^{t ^{2}} 1/\sqrt{1+u{2}} du find F"(2).\]

OpenStudy (anonymous):

i got cos(5x)^2*5+-cos(cosx)^2*-sinx

OpenStudy (anonymous):

Is this what you mean?\[f(t)=\int_1^{t^2}\frac{du}{\sqrt{1+u^2}}\]

OpenStudy (anonymous):

And are we finding \(F(t)=\int f(t)dt\) or \(F''(2)\)? Could you be a little clearer?

OpenStudy (anonymous):

i got cos(5x)^2*5+-cos(cosx)^2*-sinx

OpenStudy (anonymous):

can i send u a pdf? for one on one?

OpenStudy (anonymous):

i got cos(5x)^2*5+-cos(cosx)^2*-sinx

OpenStudy (anonymous):

ill give u medals for all the help

OpenStudy (anonymous):

Yep you can attach a file here.

OpenStudy (anonymous):

i got cos(5x)^2*5+-cos(cosx)^2*-sinx

OpenStudy (anonymous):

i got cos(5x)^2*5+-cos(cosx)^2*-sinx

OpenStudy (anonymous):

#2

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

Sorry. That integral isn't a regular trig integral.\[\int \frac{du}{\sqrt{1+u^2}}=\sinh^{-1}(u)+C\]Have you seen those before in your class?

OpenStudy (anonymous):

eh in the book perhaps, not lecture

OpenStudy (anonymous):

hmm ooh actually you don't need to know that. Notice this:\[F'(x)=\frac{d}{dx}\int_1^x f(t)dt=f(x)\]So\[F''(t)=f'(t)\]Can you figure out what \(f'(t)\) is (it's similar to the above but slightly more involved...you need the chain rule since there's a \(t^2\) in the bounds)?

OpenStudy (anonymous):

2 \[\sqrt{1+u^{4}}\]

OpenStudy (anonymous):

Very close but not quite...one little thing you need to add. Also, it should be in terms of \(t\), but that's a subtlety and won't actually affect your answer.

OpenStudy (anonymous):

2\[\sqrt{1+t^{4}}+c\]

OpenStudy (anonymous):

Actually no constant of integration since it's a definite integral. What's \(\frac{d}{dt}t^2\)?

OpenStudy (anonymous):

2..... that's why my 2 is above the radical

OpenStudy (anonymous):

That's not true.\[\frac{d}{dt}t^2=2t\]

OpenStudy (anonymous):

oops crap!

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

So what's the final answer?

OpenStudy (anonymous):

2t\[\sqrt{1+t^{4}}\]

OpenStudy (anonymous):

The question asked \(F''(2)\) not \(F''(t)\).

OpenStudy (anonymous):

lol that little stuff is going to tear me up! thanks for being so patient! usually im not this bad i promise .... 2*2= 4 ..... 2^4=16+1=17 so 4/17=final

OpenStudy (zarkon):

close

OpenStudy (anonymous):

LMFAO sqrt!!!!!! 4/\[\sqrt{17}\]

OpenStudy (anonymous):

okay okay im done being stupid putting my thinking cap

OpenStudy (anonymous):

yakeyglee want to help with #3? i got cos(5x)^2*5+-cos(cosx)^2*-sinx